# Condensed Sample-Guided Model Inversion for Knowledge Distillation

### v1

@@ Introduction
Table 2: Impact of different types of condensed samples (CS) on student accuracy (%) for different dataset condensation strategies. Condensed dataset with 50 spc from CIFAR10 was used.
@@
Table 3: MobileNet-v2 student accuracy on CIFAR10 for various configurations of the components used in our method. Condensed dataset with 50 spc from CIFAR10 was used.
Dataset CIFAR-10 CIFAR-100 ImageNet-200
Teacher ResNet-34 WRN-40-2 ResNet-34 WRN-40-2 ResNet-34 ResNet-34
Teacher acc. 95.70 94.87 78.05 75.83 71.20 71.20
Accuracy of student (S) trained with labeled data
Train w/ full real data 93.79 93.79 72.80 72.80 64.90 55.06
Train w/ cond. samples (CS) 27.24 27.24 10.69 10.69 3.50 5.60
Train w/ few real samples (RS) 27.55 27.55 4.66 4.66 1.42 1.30
Accuracy of student (S) trained with synthetic data
PRE-DFKD 83.12 83.12 66.56 61.83 54.20 47.26
PRE-DFKD* (ours w/ CS) 86.89 88.34 67.39 63.41 55.22 49.73
PRE-DFKD* (ours w/ RS) 86.77 87.47 71.63 64.37 54.95 49.68
Table 4: Student accuracies (%) obtained by enhancing PRE-DFKD with condensed and real-sample guidance (CS and RS). Condensed and real sample sets containing 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200 were used.
Dataset CIFAR-10 CIFAR100 Teacher VGG-11 VGG-11 VGG-11 VGG-11 Student VGG-11 (50% pruned) VGG-11 (75% pruned) VGG-11 (25% pruned) VGG-11 (50% pruned) Teacher acc. 92.25 92.25 71.23 71.23 Few-shot distillation accuracy FSKD 78.69 36.00 54.72 24.73 PRE-DFKD* 83.26 68.48 63.81 61.21 ((a)) Student accuracy comparison with FSKD baseline. Real sample sets containing 50 spc from CIFAR10, and 10 spc from CIFAR100 were used. Dataset CIFAR-10 CIFAR100 Teacher VGG-16 VGG-16 Student ResNet-18 ResNet-18 Teacher acc. 94.16 74.00 Few-shot distillation accuracy NetGraft 73.69 55.51 PRE-DFKD* 88.38 68.54 ((b)) Comparison with few-shot NetGraft baseline. 1 spc from both CIFAR10 / 100 were used.
This study involves student accuracies upon enhancing a base model inversion method (CMI) in 3 different settings. (1) utilizing condensed samples only by appending them to the distillation set (“+ CS”), (2) using them also to regularize the feature distribution of synthetic data (“+ CS-guided MI”) (3) the final version of our condensed sample-guided model inversion (CMI*). As seen in Table 3 “+ CS-guided MI” achieves advantage over simply including condensed samples in the distillation set (“+ CS”). Later, we include the final component, i.e., class-specific alignment, and record student accuracy (“+ class-specific alignment”). Although this addition caused only a marginal increase in performance, we still retain it as it improves results without introducing any notable computational burden.
Real-data guided sample synthesis
Our method is also inherently capable of utilizing limited real samples from the target dataset when provided. To show this, we repeated the experiments in Table 4 by replacing the condensed samples with real ones. We assume the availability of the same amount of samples randomly drawn from the target datasets as we used in experiments with condensed data (50 spc for CIFAR10, 10 spc for CIFAR100 and 10 spc for ImageNet-200). The results are displayed in Table 4 as PRE-DFKD* (ours w/ RS). Similar to condensed sample experiments, the improvements observed for heterogeneous pairs are again greater than the homogeneous ones, i.e. ResNet-34 & ResNet-18. The overall improvement yielded by our method upon utilizing few real samples is comparable that using condensed ones. This highlights the effectiveness of our approach in few-shot KD scenarios.
@@
We benchmarked the effectiveness of our method in few-shot KD against FSKD and Netgraft baselines using the same experiment setups reported in their papers. To ensure a fair evaluation of our work against these few-shot baselines, we selected identical teacher-student pairs to those in the original papers. We again use the same amount of samples from both datasets as in Table 4. In our comparison with FSKD, we use VGG11 teachers and student models that are channel-pruned versions of the teacher at different rates. As for the comparison with Netgraft, we use VGG16 teachers and ResNet-18 students, with 1 spc from both datasets. From Table 5, it can be observed that FSKD is only effective for low pruning rates while performing poorly for higher rates. Similarly, the accuracies achieved by NetGraft were also much lower than the upper-bound as seen in Table 5. This is expected as both baselines are restricted to the limited information made available by the few-shot sample set. As our method has a generative component, it does not have such restriction and therefore outperforms both few-shot baselines by a large margin.
@@ Conclusion
In conclusion, our study addresses the challenge of enhancing data-free knowledge distillation by incorporating supplementary information available from the target dataset—something traditional model inversion methods fail to leverage. Recognizing the potential of condensed samples to serve as this supplementary information, we developed a method that refines the synthetic data generation process within the model inversion framework. This integration helps align the synthetic data more closely with the real data distribution, effectively reducing the domain gap that often undermines KD performance. Our experiments demonstrate the effectiveness of this approach, showing accuracy improvements of up to $11.4\%$ across various datasets and model configurations. The method is also applicable in environments where only a few real samples are accessible, such as in few-shot learning scenarios.
Acknowledgement
This work is supported by the National Research Foundation, Singapore under its Competitive Research Programme Award NRF-CRP23-2019-0003 and Singapore Ministry of Education Academic Research Fund T1 251RES1905. We thank Cihan Acar and Liu Siying for helpful comments on the manuscript.
Appendix A: Computing Infrastructure
The specifications of the system used to run computational experiments are:
  • Ubuntu 22.04.4 LTS
  • NVIDIA A100 80GB GPUs
  • Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz
Appendix B: Implementation Details
Generator architectures
The Fast and CMI baselines share the same generator architecture, which is detailed in Table 6. The PRE-DFKD generator differs slightly, with the inclusion of spectral normalization following the convolution operations, as outlined in Table 7. In addition, the final activation function is hyperbolic tangent (Tanh), as opposed to sigmoid. Both tables utilize ”$nz$” to represent the dimension of the random vector that is drawn from a standard normal distribution. In PRE-DFKD, the value of ”$nz$” is set to 1000, while in Fast and CMI, it is set to 256. Lastly, the described architectures are cofigured to produce 32x32 RGB images (for CIFAR-10 and CIFAR-100). The activation shapes vary for datasets with samples of higher resolution (e.g. ImageNet-200).
Layer # Operation in shape out shape
1 Linear nz 8x8x128
2 BN 8x8x128 8x8x128
3 2xUpsample 8x8x128 16x16x128
4 Conv-BN-ReLU 16x16x128 16x16x128
5 2xUpsample 16x16x128 32x32x128
6 Conv-BN-ReLU 32x32x128 32x32x64
7 Conv 32x32x64 32x32x3
8 Sigmoid 32x32x3 32x32x3
Table 6: Generator architecture used in Fast and CMI experiments
Layer # Operation in shape out shape
1 Linear nz 8x8x128
2 BN 8x8x128 8x8x128
3 2xUpsample 8x8x128 16x16x128
4 Conv-SN-BN-ReLU 16x16x128 16x16x128
5 2xUpsample 16x16x128 32x32x128
6 Conv-SN-BN-ReLU 32x32x128 32x32x64
7 Conv-SN 32x32x64 32x32x3
8 Tanh 32x32x3 32x32x3
9 BN 32x32x3 32x32x3
Table 7: Generator architecture used in PRE-DFKD experiments. SN stands for spectral normalization.
Discriminator architecture & training details
For all experiments, we utilize the same discriminator whose architecture is described in Table 3. “nz” has the same meaning as in Tables 6 and 7. “nc” corresponds to the number of classes, for instance “nc=100” for CIFAR-100 dataset. To update the discriminator we use Adam optimizer with learning rate=$0.002$ and beta values=$[0.5,0.999]$. While propagating the discriminator loss to the generator, we scale it with the coefficient of $0.5$, thus the generator loss $L_{G}$ in Equation 2 becomes:
${L}_{G}=\mathcal{L}_{MI}+0.5\times\mathcal{L}_{FA}$ (6)
Layer # Operation in shape out shape
1 Linear nz+nc 1024
2 ReLU 1024 1024
3 Dropout (0.5) 1024 1024
4 Linear 1024 1
Table 8: Architectural details of the feature discriminator.
Coupling with model inversion methods
Fast and CMI use only a single generator, so we simply attach our discriminator as shown in Figure 1(a) in the main paper. In addition, we notice that both methods reset the generator periodically by randomly re-initializing its weights. To prevent any instability in our adversarial feature regularization setup, we also reset the discriminator with the same frequency. We employ the Fast-2 configuration on CIFAR-10 and CIFAR-100 datasets, which update the generator two times for every update that the student receives. Meanwhile, for ImageNet-200, we adhere to the Fast-50 configuration, which is consistent with the original paper’s approach for complex datasets. As the PRE-DFKD framework contains 2 generators, namely “novel” and “memory” generators, attaching our method requires further consideration. While the “novel” generator is responsible for producing new/informative samples in the framework, the memory generator memorizes the cumulative output distribution of the novel generator. As conditioning the novel generator also implicitly affects the memory generator, we only pair the discriminator with it. We run CMI, Fast, and PRE-DFKD for 200, 220, and 220 epochs, respectively. All results reflect the maximum student accuracy observed during distillation.
Moreover, PRE-DFKD does not store generated samples and instead distills the student on-the-fly with newly generated samples at each epoch. This prevents us from directly applying Algorithm 1 to combine condensed and synthetic samples, and the other alternatives discussed in Section 3 impair student accuracy. For this reason, we only use the condensed samples to guide synthetic data generation and exclude them while training the student. Even without utilizing condensed samples in a direct way in distillation, we observe accuracy improvement on both datasets.
Appendix C: Condensed Datasets
The basic dataset condensation process uses a reference model and optimizes condensed samples to match the gradients that larger real batches produce for that model. In our experiments, we used condensed samples retrieved from the official GitHub pages of (Zhao, Mopuri, and Bilen 2021) and (Cazenavette et al. 2022). Both works, used the same 3-layer CNN architecture as the reference model to generate these datasets. For CIFAR-10, we used samples generated by both (Zhao and Bilen 2021) and (Zhao and Bilen 2023). The samples from (Zhao and Bilen 2021) are available in three sizes: 1, 10, and 50 samples per class (spc), while (Zhao and Bilen 2023) provides samples only at 50 spc. For CIFAR-100, only the 1 and 10 spc settings from (Zhao and Bilen 2021) are available. For ImageNet-200, we utilized the condensed samples from (Cazenavette et al. 2022), which provides 10 spc. It is worth noting that none of the teacher models in our experiments were used as reference models during the generation of these samples. This demonstrates that our method can effectively utilize condensed datasets, regardless of the reference model used in their optimization process.
Appendix D: Time complexity of model inversion
To show that our method does not sacrifice significant speed while enhancing the performance, we share the model inversion times in table 9.
Method Time (hrs)
Fast 3
Fast* 4
CMI 11
CMI* 13.5
PRE-DFKD 11.6
PRE-DFKD* 13.5
Table 9: Times consumed by different model inversion methods.
Appendix E: Class-specific Feature Alignment
When the discriminator is configured to only differentiate between the generated synthetic samples and the condensed ones, the generator can achieve its optimization objective by merely aligning the two distributions, as demonstrated in Figure 6(a). However, for optimal results, it is preferable to align the synthetic samples with the condensed ones based on their respective classes, as shown in Figure 6(b). For instance, a synthetic car image should exhibit similar semantic features to a condensed car image rather than a condensed sunflower image.
((a)) generic alignment
((b)) class-specific alignment
Figure 6: Illustration of feature spaces observed after different feature alignment strategies. Condensed (C) and synthetic (S) features are denoted by red and blue colors respectively. Each shape indicates a unique class ID (e.g. c1).

### v2

@@ Introduction
Table 2: Impact of different types of condensed samples (CS) on student accuracy (%) for different dataset condensation strategies. Dataset contains 50 spc from CIFAR10.
DSA, DM and MTT refer to (zhao2021siamese), (zhao2023dataset), and (cazenavette2022dataset) respectively. The results indicate that while the choice of condensation method can impact the final student accuracy, our method consistently improves performance across all types of condensed samples tested. This suggests that its effectiveness is not dependent on any single condensation approach.
Dataset CIFAR-10 CIFAR-100 ImageNet-200
Teacher ResNet-34 WRN-40-2 ResNet-34 WRN-40-2 ResNet-34 ResNet-34
Teacher acc. 95.70 94.87 78.05 75.83 71.20 71.20
[…]
Train w/ full real data 93.79 93.79 72.80 72.80 64.90 55.06
Train w/ cond. samples (CS) 27.24 27.24 10.69 10.69 3.50 5.60
Train w/ few real samples (RS) 27.55 27.55 4.66 4.66 1.42 1.30
[…]
PRE-DFKD 83.12 83.12 66.56 61.83 54.20 47.26
PRE-DFKD* (ours w/ CS) 86.89 88.34 67.39 63.41 55.22 49.73
PRE-DFKD* (ours w/ RS) 86.77 87.47 71.63 64.37 54.95 49.68
Table 3: Student accuracies (%) obtained by utilizing the expanded condensed and real datasets set using PRE-DFKD (CS and RS). Condensed and real sample sets contain 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200.
[…]
@@
Table 5: MobileNet-v2 student accuracy on CIFAR10 for various configurations of the components used in our method. Dataset with 50 spc from CIFAR10 was used.
This involves student accuracies upon enhancing a base model inversion method (CMI) in 3 different settings. (1) utilizing condensed samples only by appending them to the distillation set (“+ CS”), (2) using them also to regularize the feature distribution of synthetic data (“+ CS-guided MI”) (3) the final version of our condensed sample-guided model inversion (CMI*). As seen in Table 5 “+ CS-guided MI” achieves advantage over simply including condensed samples in the distillation set (“+ CS”). Later, we include the final component, i.e., class-specific alignment, and record student accuracy (“+ class-specific alignment”). This caused only a marginal increase in performance but we retain it as it improves results without introducing any notable computational burden.
Impact of expanded real data on KD
Our method is also inherently capable of expanding limited real samples from the target dataset in few-shot scenarios. To show this, we repeated the experiments in Table 3 by replacing the condensed samples with real ones. We assume the availability of the same amount of samples randomly drawn from the target datasets as we used in experiments with condensed data (50 spc for CIFAR10, 10 spc for CIFAR100 and 10 spc for ImageNet-200). The results are displayed in Table 3 as PRE-DFKD* (ours w/ RS). Similar to condensed sample experiments, the improvements observed for heterogeneous pairs are again greater than the homogeneous ones, i.e. ResNet-34 & ResNet-18. The overall improvement yielded by our method upon utilizing few real samples is comparable that using condensed ones.
@@
We benchmarked the effectiveness of our method in few-shot KD against FSKD and Netgraft baselines using the same experiment setups reported in their papers. To ensure a fair evaluation of our work against these few-shot baselines, we selected identical teacher-student pairs to those in the original papers. We again use the same amount of samples from both datasets as in Table 3. In our comparison with FSKD, we use VGG11 teachers and student models that are channel-pruned versions of the teacher at different rates. As for the comparison with Netgraft, we use VGG16 teachers and ResNet-18 students, with 1 spc from both datasets. From Table 4, it can be observed FSKD is only effective for low pruning rates while performing poorly for higher rates. Similarly, the accuracies achieved by NetGraft were also much lower than the upper-bound as seen in Table 4. This is expected as both baselines are restricted to the limited information made available by the few-shot sample set. As our method has a generative component, it does not have such restriction and therefore outperforms both few-shot baselines by a large margin.
@@ Conclusion
We address the challenge of limited information in condensed datasets, which often hinders their effectiveness in KD. We propose a method that leverages MI to expand these condensed datasets, generating synthetic data that enriches the compact representation and improves KD performance. Our approach utilizes condensed samples as prototypes to guide the MI process, ensuring that the generated synthetic data aligns closely with the underlying data distribution represented by the condensed set. Our experiments demonstrate the effectiveness of our method across various datasets, model architectures, and state-of-the-art MI techniques (Fast, CMI, PRE-DFKD). The results show consistent improvements in KD, compared to using condensed datasets alone or standard MI-based KD methods. Our approach also applies to few-shot learning, outperforming existing few-shot KD by effectively leveraging limited real data for synthetic data generation.